OpenStudy (anonymous):

let f(x)=x^2-13 and g(x)=16-x. Perform the composition or operation indicated. (f\g)(-9)

OpenStudy (tanvirzawad):

Did you ask what will be the value if I plugin -9 into (x^2-13)/(16-x)? In that case the answer will be 2.72. - - - Try this: let, h(x)=(16-x)^2-13. Now find h(-9). [note that (a-b)^2=a^2-2ab+b^2]

hero (hero):

(f/g)(x) = f(x)/g(x) (f/g)(-9) = (-9)^2 - 13/(16+9) (f/g)(-9) = 81 - 13/25 (f/g)(-3) = 68/25